# I. Márquez 11    C. Boisson 22    M. Joly 22    D. Pelat 22    F. Durret 33

### v1

@@ 4 Spectral fits
The velocity dispersions, $\sigma$, of the galaxies observed with ISAAC, were measured by cross-correlation of the full observed H-band spectrum of the galaxy with infrared template stars (see Boisson et al., 2000, 2002). The resulting $\sigma$ values are given in Tables 3 to 9. The differential velocity dispersion between galaxies and template stellar spectra is handled in the synthesis.
@@ 4.1 MCG-6-30-15
Figure 1: Sharp divided optical HST image in the F547M filter of MCG-6-30-15. The region studied by Raimundo et al. (2013) that we call ”pseudo-ring” corresponds to the large blue circle (1.1 arcsec radius) minus the small blue rectangle, and their regions labeled 1 to 4 are shown in red. The ISAAC slit is superimposed in green, with the centre and the two regions shown as the central and external rectangles respectively.
@@
Figure 3: Stellar population fit to the ”pseudo-ring” of MCG-6-30-15 (see Fig. 1 top).
Figure 4: Stellar population fits to the three regions of MCG-6-30-15 taken with VLT/ISAAC. From top to bottom: nucleus, East and West (see Fig. 1 bottom)
@@
The high spatial resolution of SINFONI allows to well sample the inner nuclear regions. We analyse the five SINFONI regions (1 to 4 and a pseudo-ring of radius $\sim$1.1 arcsec excluding the very nucleus) observed by Raimundo et al. (2013). The ISAAC spectra sample two more regions besides the ones mapped by SINFONI: the nucleus ($<$ 1 arcsec /110 pc radius) and two regions east and west. The regions studied are marked on the sharp-divided image of the central region of MCG-6-30-15 as shown in Fig. 1.
The results of the stellar population of the various regions of MCG-6-30-15 are given in Table 3, with the corresponding spectra and fits shown in Figs. 2 and 3. Note the excellent agreement between the observed and synthetic spectra.
@@
The whole central $<$110 pc radius population is old ($\sim$ 5.0 $\times$109 yrs), dominated by dwarfs and evolved M giant stars. An important contribution of young (107 yrs) supergiant stars (M2I) is also present. Using emission lines to study the supernova rate, Raimundo et al. (2017) also deduce the presence of a young $<$108 yrs population. Note that the better spectral resolution of SINFONI, compared to ISAAC, reveals the presence of old super metal-rich dwarf stars as expected in case of successive bursts of star formation.
As can be seen from the ISAAC data, the outer regions (110-330 pc) are dominated by intermediate (F2V, $\sim$ 4.0 $\times$108 yrs) and old stars with no recent burst of star formation as indeed expected in lenticular galaxies (see Fig. 4). At such a large distance from the nucleus, the metallicity of the contributing stars is weaker than in the nuclear regions. A  20% contribution of young star population is present in the nucleus ( $<$ 110 pc).
Analysing in a large IR range (J, H and K bands) the inner 300 pc, Riffel et al. (2024) found a dominating old population, the youngest contribution they obtain is of intermediate age ($>$108 yrs). This may be due to a lack of spatial resolution.
In conclusion, the nuclear stellar content of the Seyfert 1 MCG–6-30-15 is dominated by a mixture of old and young populations with the contribution of metal-rich stars proving a history of multiple bursts of star formation. This is in contrast with the overall bulge population inferred from our study in the optical and points to a population gradient with ongoing star formation in the nucleus.
@@ 4.2 NGC 2110
Figure 5: Sharp divided infrared HST image in the fr680p15 filter of NGC 2110 with the ISAAC slit superimposed and the studied regions marked.
Figure 6: Stellar population fits to the four regions of NGC 2110. From top to bottom: the nucleus, the SE and the two NW regions (see Fig. 5)
@@
The stellar population in the nucleus of NGC 2110 is dominated by giant evolved stars. In the ring surrounding the nucleus, up to 120 pc, an old population is still dominating, but includes partly metal-rich and partly weak metallicity dwarf stars. The farthest NW region, beyond 120 pc of the centre, shows a large contribution of intermediate stars indicative of an ancient ($\sim$ 3.0 $\times$108 yrs) star formation burst in the past history of the galaxy. This population gradient is in line with the results of a study of the inner 3x3 arcsec2 (450x450 pc2) by Diniz et al. (2019) through near infrared J and K bands integral field spectroscopy.
@@ 4.3 NGC 2992
Figure 7: Sharp divided infrared HST image in the F196N filter of NGC 2992 with the ISAAC slit superimposed and the three regions indicated: the nucleus (C), the north-east (NE) and south-west (SW) regions.
Figure 8: Stellar population fits to the three regions of NGC 2992. From top to bottom: nucleus, SW and NE (see Fig. 7).
@@
In a detailed IR adaptive optics study, Chapman et al. (2000) found a radial distribution of the CO index in the inner 3 arcsec ($\sim$ 450 pc), suggesting a population gradient within the core, with the stellar population at the very centre being older than that in the surrounding regions. Stellar population modelling of OASIS integral field spectroscopy by Stoklasová et al. (2009) revealed a contribution of young stars close to the dust lane, younger than in the rest of their field of view, with mostly old stars in the eastern part. From optical GMOS observations, Guolo-Pereira et al. (2021) find a mainly old metal-rich stellar population, with a noticeable contribution of young metal-poor stars in the inner 1.1 kpc of NGC 2992. Burtscher et al. (2021) from VLT/X-shooter spectra in an aperture of 150 parsec radius infer a dominating rather old reddened stellar population ($>10^{9}$ yrs). Another study by Dahmer-Hahn et al. (2022) with MUSE integral field spectroscopy of the 270 central parsecs, points to a mainly young, metal-rich and highly reddened stellar population to the north–west, the rest of their FoV having an old, metal-poor and less reddened stellar population. However a study in the IR by Riffel et al. (2024), with SpeX attached to the IRTF, sampling the same dimensions, infers a largely dominating old population ($\sim$ 8 Gyr).
In Table 5 the results of the stellar population fits for the three regions of NGC 2992 we studied are given, with the corresponding spectra and fits shown in Fig. 8.
The AGN non stellar contribution reaches about 60% in the nucleus. We find an evolved (M giant stars, $4\times 10^{9}$ yrs) stellar population of about solar metallicity, in a high extinction environment, with no hint of a young stellar population contribution. The stellar population is younger in the regions surrounding the nucleus, with not more than the foreground galactic extinction. An important star formation difference is visible between the SW and NE regions: presence of an important metal-rich component and recent ($\sim 10^{7}$ yrs) starburst in the SW region, the closest to the dust lane, and an older burst of star formation inducing an intermediate age stellar population in the NE region.
The hot dust temperature inferred by the analysis of the SW region is not expected. A test constraining the dust temperature to 500 K induces an increase of the computed extinction up to the nuclear one, a plausible result as this region is the closest to the dust lane. The global stellar population remains the same, pointing to some degeneracy between extinction and featureless continuum from dust emission.
In our study of the 600 central parsecs, taking advantage of the H-band, we can emphasise a clear age and possible metallicity gradient, besides a high star formation in the SW region.
@@ 4.4 NGC 3185
Figure 9: Sharp divided HST image in the F814W filter of NGC 3185 with the ISAAC slit superimposed. The nuclear region is indicated by C and the ring corresponds to the sum of the two blue regions.
@@
NGC 3185 is a strongly-barred SBa galaxy located in a compact Hickson group dominated by NGC 3190. This galaxy classified as Seyfert 2 by Filippenko & Sargent (1985), is also considered as a transition object with both Seyfert and star forming characteristics (e.g., Gonçalves et al., 1999). A LINER classification was also proposed by Carrillo et al. (1999). Circumgalactic rings of star-forming material are observed in both UV and IR images on a large scale (Lanz et al., 2013).
@@
Contrary to what is expected for a genuine Seyfert nucleus, no dilution by a power law component is needed to account for the nuclear spectrum. Within the central 360 pc of NGC 3185, the stellar population is old, partly metal-rich, though intermediate-age dwarf F stars in the ring indicate some stellar activity in the past. No extinction is found in excess of the galactic value. This points toward a currently quiet stellar environment and low activity in the central engine.
@@ 4.5 NGC 3783
Figure 11: HST image in the F160W filter of NGC 3783 with the ISAAC slit superimposed. The ring corresponds to the sum of the two regions labeled R.
@@
NGC 3783 is a nearly face-on SBa galaxy with an unobscured very bright and variable Seyfert 1 nucleus (e.g., Bentz et al., 2009). The sharp-divided image of NGC 3783 is shown in Fig. 11 with the studied region superimposed ; it clearly shows the high contamination from the Sy1 nucleus in this galaxy.
@@
We made no attempt to recover the nuclear stellar population from our data in the central 140 pc, since the very strong AGN emission overwhelms the nuclear flux. As NGC 3783 is viewed almost face-on with a faint symmetrical bulge structure, a hereafter called ring-like spectrum is built by adding the very similar spectra extracted on either side of the nucleus (Boisson et al., 2002). The inferred stellar population is given in Table 7, and the corresponding spectrum and the fit are shown in Fig. 12.
Though the stellar population is dominated by about 65% of old and evolved stars ($>$ 4.0 $\times$ 109yrs), a strong contribution of intermediate age stars (33%) is present as well as a hint of a recent star formation event. We may see the trace of a young population in the surrounding region of NGC 3783.
@@ 4.6 NGC 6221
Figure 13: Sharp divided infrared HST image in the F160N filter of NGC 6221 with the ISAAC slit superimposed. The three regions are the nucleus (C) and the two regions to the north, N1 and N2.
@@
NGC 6221, is a barred Sbc galaxy located within a small galaxy group, with a bar clearly visible in the optical and infrared. This X-ray bright Seyfert galaxy is known to exhibit clear signs of intense star formation together with a weak Seyfert nucleus, exhibiting both signatures of a Seyfert 1 in the X-rays and a typical reddened (Av=3) starburst-like spectrum in the optical (e.g., Levenson et al., 2001; La Franca et al., 2016).
The sharp-divided image of NGC 6221 is shown in Fig. 13 with the studied regions superimposed. We study the nuclear spectrum within 80 pc, and two regions towards the north up to 200 pc, excluding from our study the south conspicuous HII region spectrum (Boisson et al., 2002). The corresponding stellar populations are given in Table 8, with spectra and fits shown in Fig. 14. Note that the Brackett (4-13) emission line at 1.61 $\mu$m visible on all spectra has been excluded from the fit.
@@ 4.7 NGC 7582
Figure 15: Sharp divided infrared HST image in the F606WN filter of NGC 7582 with the ISAAC slit superimposed. The north (N1, N2) and south (S1, S2) regions are indicated.
Figure 16: Stellar population fits to the four regions of NGC 7582. From top to bottom: N2, N1, S1 and S2 (see Fig. 15).
NGC 7582 is a nearby barred galaxy in the Grus quartet that harbours a nuclear starburst surrounding the active nucleus (Wold & Galliano, 2006; Bianchi et al., 2007). The Galactic column density toward NGC 7582 is very low. This X-ray bright Seyfert 2 galaxy (Bianchi et al., 2009) is a typical “changing look AGN”, a broad emission line component being reported by Aretxaga et al. (1999). This component was initially proposed as caused by a supernova explosion, but was later confirmed to be from an AGN (Ricci et al., 2018).
The sharp-divided image of NGC 7582 (Fig. 15) shows a disturbed morphology with an obscured nucleus, and a dust lane crossing over the nuclear region. Due to strong dilution by the AGN non-thermal continuum, the nuclear spectrum (80 arcsec diameter) is useless for stellar population studies. However two external regions on each side of the nucleus were extracted, whose stellar populations are given in Table 9, with the corresponding spectra and fits shown in Fig. 16. The northern regions are shining through dust patches. This translates into higher E(B-V) in those regions than on the other side of the nucleus. Note that the Brackett (4-13) emission line at 1.61 $\mu$m visible on all spectra has been excluded from the fit.
We find an important population of evolved giant stars ($\sim(4-6)\times 10^{9}$ yrs), intermediate ones ($\sim 4\times 10^{8}$ yrs) as well as very young (107 yrs) stars in the four selected regions. This is in line with what is found in the optical range by Burtscher et al. (2021) and in the near IR by Riffel et al. (2024) within the 200 central parsecs.
@@ 5 Summary and conclusions
Taking advantage of the rich near-IR H-band, we investigated the stellar populations of the central few 100  pc of the host galaxy of seven active galactic nuclei, with the help of a stellar spectral database allowing to detect the young star contribution down to $\leq 10^{7}$ yrs.
Our results can be summarized as follows:
First, in the inner 100 pc, we detect some metal-rich population in all AGN but NGC 7582. The only Seyfert 1 for which we could synthesize in detail the nucleus, MCG 06-30-15, shows a strong nuclear starburst. The two composite Seyfert 2s, NGC 6221 and NGC 7582, are also dominated by a starburst component, while no young stars are detected in the genuine Seyfert 2 galaxies, NGC 2110, and NGC 3185, nor in NGC 2992.
Moreover, with our good spatial sampling, we could show how the star formation history of the inner nucleus of the galaxies is highly heterogeneous, as young bursts of star formation, intermediate (poststarburst) and old stellar populations all appear in widely varying proportions at all places. For example, one can note that in the four regions within 100 pc of MCG 06-30-15 (for which we have the best spatial and spectral resolutions), we find a large variety of populations. We could highlight such a diversity thanks to the infrared domain, rich in stellar features and where the star contribution is higher. There thus may not be only one process / relation between fuelling by star formation and power of the AGN.
@@
These results indicate that an active nucleus always seems to be linked to some past or even ongoing stellar formation. This was also noted by e.g. Davies et al. (2007) or Riffel et al. (2022) who, based on high spatial resolution integral field observations, found that the near-IR light in the inner few 100 pc of active galaxies is primarily composed on average of moderately young (a few $10^{7}$ to $10^{8}$ yrs) to intermediate-age (a few $10^{8}$ yrs) stellar populations, and that the age of the starburst seems connected to the AGN’s luminosity. However those studies concern only the post-starburst population, with ages $\sim 10^{8}$$10^{9}$ yr.
Based on results of the stellar population study, in the optical, of a sample of 12 galaxies covering a range of morphology types and nuclear activities, Boisson et al. (2000) suggested the existence of a relationship between stellar population and activity type. The picture emerging from this study was that of an evolutionary sequence rather than a strict unified scheme for AGN. Indeed, differences in the stellar populations within the nuclear regions were found to follow the degree of nuclear activity. Seyfert 2 galaxies appear to exhibit fossil star formation activity, i.e. a stellar population younger than found in normal galaxies, but somewhat older and more metal-rich than found in starburst galaxies. In an optical study of the population synthesis of the inner 300 pc of star-forming and active galaxies, including the three Seyfert 2 galaxies of our sample, Burtscher et al. (2021) found a dominance of young stellar populations too, while this was not the case for their control sample of inactive galaxies. They also showed that such young populations are not necessarily indicative of ongoing star formation. This is along the lines of the evolutionary sequence of Boisson et al. (2000, 2004). Based on the similar nuclear star formation histories of the Seyfert 2s and the star-forming control sample galaxies, Burtscher et al. (2021) further speculate that the starbursts probably turned into Seyferts for some fraction of their time.
In the present paper, in the IR, we find that indeed star-forming galaxies may turn into Seyfert 1s, while star formation in Seyfert 2s, if any, appears outside the nuclear region. Note that NGC 2992, for example, is known to change from type 1 to type 2.

### v2

@@ 4 Spectral fits
The velocity dispersions, $\sigma$, of the galaxies observed with ISAAC, were measured by cross-correlation of the full observed H-band spectrum of the galaxy with IR template stars (see Boisson et al., 2000, 2002). The resulting $\sigma$ values are given in Tables 3 to 9. The differential velocity dispersion between galaxies and template stellar spectra is handled in the synthesis.
@@ 4.1 MCG-6-30-15
Figure 1: Sharp-divided optical HST image in the F547M filter of MCG-6-30-15. The region studied by Raimundo et al. (2013) that we call ’pseudo-ring’ corresponds to the large blue circle (1.1 arcsec radius) minus the small blue rectangle, and their regions labelled 1 to 4 are shown in red. The ISAAC slit is superimposed in green, with the centre and the two regions shown as the central and external rectangles respectively.
@@
Figure 3: Stellar population fit to the ’pseudo-ring’ of MCG-6-30-15 (see Fig. 1 top).
Figure 4: Stellar population fits to the three regions of MCG-6-30-15 taken with VLT/ISAAC. From top to bottom: nucleus, east, and west (see Fig. 1 bottom)
@@
With its high spatial resolution, SINFONI allows one to sample the inner nuclear regions well. We analyse the five SINFONI regions (1 to 4 and a pseudo-ring of radius $\sim$1.1 arcsec excluding the very nucleus) observed by Raimundo et al. (2013). The ISAAC spectra sample two more regions besides the ones mapped by SINFONI: the nucleus ($<$ 1 arcsec /110 pc radius) and two regions east and west. The regions studied are marked on the sharp-divided image of the central region of MCG-6-30-15 as shown in Fig. 1.
The results of the stellar population fits of the various regions of MCG-6-30-15 are given in Table 3, with the corresponding spectra and fits shown in Figs. 2 and 3. We note the excellent agreement between the observed and synthetic spectra.
@@
The whole central $<$110 pc radius population is old ($\sim$ 5.0 $\times$109 yrs), dominated by dwarfs and evolved M giant stars. An important contribution of young (107 yrs) supergiant stars (M2I) is also present. Using emission lines to study the supernova rate, Raimundo et al. (2017) also deduce the presence of a young $<$108 yrs population. We note that the better spectral resolution of SINFONI, compared to ISAAC, reveals the presence of old super metal-rich dwarf stars as expected in case of successive bursts of star formation.
As can be seen from the ISAAC data, the outer regions (110-330 pc) are dominated by intermediate (F2V, $\sim$ 4.0 $\times$108 yrs) and old stars with no recent burst of star formation as indeed expected in lenticular galaxies (see Fig. 4). At such a large distance from the nucleus, the metallicity of the contributing stars is weaker than in the nuclear regions. A $\sim$ 20% contribution of young star population is present in the nucleus ( $<$ 110 pc).
Analysing in a large IR range (J, H, and K bands) the inner 300 pc, Riffel et al. (2024) found a dominating old population, the youngest contribution they obtain is of intermediate age ($>$108 yrs). This may be due to a lack of spatial resolution.
In conclusion, the nuclear stellar content of the Seyfert 1 MCG–6-30-15 is dominated by a mixture of old and young populations with the contribution of metal-rich stars proving a history of multiple bursts of star formation. This is in contrast with the overall bulge population inferred from our study in the optical, and points to a population gradient with ongoing star formation in the nucleus.
@@ 4.2 NGC 2110
Figure 5: Sharp-divided IR HST image in the fr680p15 filter of NGC 2110 with the ISAAC slit superimposed and the studied regions marked.
Figure 6: Stellar population fits to the four regions of NGC 2110. From top to bottom: the nucleus, the SE, and the two NW regions (see Fig. 5)
@@
The stellar population in the nucleus of NGC 2110 is dominated by giant evolved stars. In the ring surrounding the nucleus, up to 120 pc, an old population is still dominating, but includes partly metal-rich and partly weak metallicity dwarf stars. The farthest NW region, beyond 120 pc of the centre, shows a large contribution of intermediate stars indicative of an ancient ($\sim$ 3.0 $\times$108 yrs) star formation burst in the past history of the galaxy. This population gradient is in line with the results of a study of the inner 3x3 arcsec2 (450x450 pc2) by Diniz et al. (2019) through near IR J and K bands integral field spectroscopy.
@@ 4.3 NGC 2992
Figure 7: Sharp-divided IR HST image in the F196N filter of NGC 2992 with the ISAAC slit superimposed and the three regions indicated: the nucleus (C), the north-east (NE), and south-west (SW) regions.
Figure 8: Stellar population fits to the three regions of NGC 2992. From top to bottom: nucleus, SW, and NE (see Fig. 7).
@@
In a detailed IR adaptive optics study, Chapman et al. (2000) found a radial distribution of the CO index in the inner 3 arcsec ($\sim$ 450 pc), suggesting a population gradient within the core, with the stellar population at the very centre being older than that in the surrounding regions. Stellar population modelling of OASIS integral field spectroscopy by Stoklasová et al. (2009) revealed a contribution of young stars close to the dust lane, younger than in the rest of their field of view, with mostly old stars in the eastern part. From optical GMOS observations, Guolo-Pereira et al. (2021) find a mainly old metal-rich stellar population, with a noticeable contribution of young metal-poor stars in the inner 1.1 kpc of NGC 2992. Burtscher et al. (2021) from VLT/X-shooter spectra in an aperture of 150 parsec radius infer a dominating rather old reddened stellar population ($>10^{9}$ yrs). Another study by Dahmer-Hahn et al. (2022) with MUSE integral field spectroscopy of the 270 central parsecs, points to a mainly young, metal-rich, and highly reddened stellar population to the north–west, with the rest of their field of view (FoV) having an old, metal-poor, and less reddened stellar population. However, a study in the IR by Riffel et al. (2024), with SpeX attached to the IRTF, sampling the same dimensions, infers a largely dominating old population ($\sim$ 8 Gyr).
In Table 5 the results of the stellar population fits for the three regions of NGC 2992 we studied are given, with the corresponding spectra and fits shown in Fig. 8. The AGN non stellar contribution reaches about 60% in the nucleus. We find an evolved (M giant stars, $4\times 10^{9}$ yrs) stellar population of about solar metallicity, in a high extinction environment, with no hint of a young stellar population contribution. The stellar population is younger in the regions surrounding the nucleus, with not more than the foreground galactic extinction. An important star formation difference is visible between the SW and NE regions: presence of an important metal-rich component and recent ($\sim 10^{7}$ yrs) starburst in the SW region, the closest to the dust lane, and an older burst of star formation inducing an intermediate age stellar population in the NE region.
The hot dust temperature inferred by the analysis of the SW region is not expected. A test constraining the dust temperature to 500 K induces an increase of the computed extinction up to the nuclear one, a plausible result as this region is the closest to the dust lane. The global stellar population remains the same, pointing to some degeneracy between extinction and featureless continuum from dust emission. In our study of the 600 central parsecs, taking advantage of the H-band, we can emphasise a clear age and possible metallicity gradient, besides a high star formation in the SW region.
@@ 4.4 NGC 3185
Figure 9: Sharp-divided HST image in the F814W filter of NGC 3185 with the ISAAC slit superimposed. The nuclear region is indicated by C, and the ring corresponds to the sum of the two blue regions.
@@
NGC 3185 is a strongly barred SBa galaxy located in a compact Hickson group dominated by NGC 3190. This galaxy classified as Seyfert 2 by Filippenko & Sargent (1985), is also considered as a transition object with both Seyfert and star forming characteristics (e.g. Gonçalves et al., 1999). A LINER classification was also proposed by Carrillo et al. (1999). Circumgalactic rings of star forming material are observed in both UV and IR images on a large scale (Lanz et al., 2013).
@@
Contrary to what is expected for a genuine Seyfert nucleus, no dilution by a power law component is needed to account for the nuclear spectrum. Within the central 360 pc of NGC 3185, the stellar population is old, partly metal-rich, though intermediate-age dwarf F stars in the ring indicate some stellar activity in the past. No extinction is found in excess of the galactic value. This points towards a currently quiet stellar environment and low activity in the central engine.
@@ 4.5 NGC 3783
Figure 11: HST image in the F160W filter of NGC 3783 with the ISAAC slit superimposed. The ring corresponds to the sum of the two regions labelled R.
@@
NGC 3783 is a nearly face-on SBa galaxy with an unobscured very bright and variable Seyfert 1 nucleus (e.g. Bentz et al., 2009). The sharp-divided image of NGC 3783 is shown in Fig. 11 with the studied region superimposed; it clearly shows the high contamination from the Sy1 nucleus in this galaxy.
@@
We made no attempt to recover the nuclear stellar population from our data in the central 140 pc, since the very strong AGN emission overwhelms the nuclear flux. As NGC 3783 is viewed almost face-on with a faint symmetrical bulge structure, a hereafter called ring-like spectrum is built by adding the very similar spectra extracted on either side of the nucleus (Boisson et al., 2002). The inferred stellar population is given in Table 7, and the corresponding spectrum and fit are shown in Fig. 12.
Though the stellar population is dominated by about 65% of old and evolved stars ($>$ 4.0 $\times$ 109yrs), a strong contribution of intermediate age stars (33%) is present as well as a hint of a recent star formation event. We may see the trace of a young population in this region surrounding the nucleus of NGC 3783.
@@ 4.6 NGC 6221
Figure 13: Sharp-divided IR HST image in the F160N filter of NGC 6221 with the ISAAC slit superimposed. The three regions are the nucleus (C) and the two regions to the north, N1 and N2.
@@
NGC 6221, is a barred Sbc galaxy located within a small galaxy group, with a bar clearly visible in the optical and IR. This X-ray bright Seyfert galaxy is known to exhibit clear signs of intense star formation together with a weak Seyfert nucleus, exhibiting both signatures of a Seyfert 1 in the X-rays and a typical reddened (Av=3) starburst-like spectrum in the optical (e.g. Levenson et al., 2001; La Franca et al., 2016).
The sharp-divided image of NGC 6221 is shown in Fig. 13 with the studied regions superimposed. We study the nuclear spectrum within 80 pc, and two regions towards the north up to 200 pc, excluding from our study the south conspicuous HII region spectrum (Boisson et al., 2002). The corresponding stellar populations are given in Table 8, with spectra and fits shown in Fig. 14. We note that the Brackett (4-13) emission line at 1.61 $\mu$m visible on all spectra has been excluded from the fit.
@@ 4.7 NGC 7582
Figure 15: Sharp-divided IR HST image in the F606WN filter of NGC 7582 with the ISAAC slit superimposed. The north (N1, N2) and south (S1, S2) regions are indicated.
Figure 16: Stellar population fits to the four regions of NGC 7582. From top to bottom: N2, N1, S1, and S2 (see Fig. 15).
NGC 7582 is a nearby barred galaxy in the Grus quartet that harbours a nuclear starburst surrounding the active nucleus (Wold & Galliano, 2006; Bianchi et al., 2007). The Galactic column density towards NGC 7582 is very low. This X-ray bright Seyfert 2 galaxy (Bianchi et al., 2009) is a typical “changing look AGN”, a broad emission line component being reported by Aretxaga et al. (1999). This component was initially proposed as caused by a supernova explosion, but was later confirmed to be from an AGN (Ricci et al., 2018).
The sharp-divided image of NGC 7582 (Fig. 15) shows a disturbed morphology with an obscured nucleus, and a dust lane crossing over the nuclear region. Due to strong dilution by the AGN non-thermal continuum, the nuclear spectrum (80 arcsec diameter) is useless for stellar population studies. However, two external regions on each side of the nucleus were extracted, whose stellar populations are given in Table 9, with the corresponding spectra and fits shown in Fig. 16. The northern regions are shining through dust patches. This translates into higher E(B-V) in those regions than on the other side of the nucleus. We note that the Brackett (4-13) emission line at 1.61 $\mu$m visible on all spectra has been excluded from the fit.
We find an important population of evolved giant stars ($\sim(4-6)\times 10^{9}$ yrs), intermediate ones ($\sim 4\times 10^{8}$ yrs) as well as very young (107 yrs) stars in the four selected regions. This is in line with what is found in the optical range by Burtscher et al. (2021), and in the near IR by Riffel et al. (2024) within the 200 central parsecs.
@@ 5 Summary and conclusions
Taking advantage of the rich near-IR H-band, we investigated the stellar populations of the central few 100  pc of the host galaxy of seven active galactic nuclei, with the help of a stellar spectral database that allowed us to detect the young star contribution down to $\leq 10^{7}$ yrs.
Our results can be summarised as follows:
First, in the inner 100 pc, we detect some metal-rich population in all AGNs but NGC 7582. The only Seyfert 1 for which we could synthesise in detail the nucleus, MCG 06-30-15, shows a strong nuclear starburst. The two composite Seyfert 2s, NGC 6221 and NGC 7582, are also dominated by a starburst component, while no young stars are detected in the genuine Seyfert 2 galaxies, NGC 2110, and NGC 3185, nor in NGC 2992.
Moreover, with our good spatial sampling, we could show how the star formation history of the inner nucleus of the galaxies is highly heterogeneous, as young bursts of star formation, intermediate (poststarburst) and old stellar populations all appear in widely varying proportions at all places. For example, one can note that in the four regions within 100 pc of MCG 06-30-15 (for which we have the best spatial and spectral resolutions), we find a large variety of populations. We could highlight such a diversity thanks to the IR domain, rich in stellar features, and where the star contribution is higher. Thus, there may not be only one process or relation between fuelling by star formation and power of the AGN.
@@
These results indicate that an active nucleus always seems to be linked to some past or even ongoing stellar formation. This was also noted by, for example, Davies et al. (2007) or Riffel et al. (2022) who, based on high spatial resolution integral field observations, found that the near-IR light in the inner few 100 pc of active galaxies is primarily composed on average of moderately young (a few $10^{7}$ to $10^{8}$ yrs) to intermediate-age (a few $10^{8}$ yrs) stellar populations, and that the age of the starburst seems connected to the AGN’s luminosity. However, those studies concern only the post-starburst population, with ages $\sim 10^{8}$$10^{9}$ yr.
Based on results of the stellar population study, in the optical, of a sample of 12 galaxies covering a range of morphology types and nuclear activities, Boisson et al. (2000) suggested the existence of a relationship between stellar population and activity type. The picture emerging from this study was that of an evolutionary sequence rather than a strict unified scheme for AGNs. Indeed, differences in the stellar populations within the nuclear regions were found to follow the degree of nuclear activity. Seyfert 2 galaxies appear to exhibit fossil star formation activity, that is, a stellar population younger than found in normal galaxies, but somewhat older and more metal-rich than found in starburst galaxies. In an optical study of the population synthesis of the inner 300 pc of star forming and active galaxies, including the three Seyfert 2 galaxies of our sample, Burtscher et al. (2021) found a dominance of young stellar populations too, while this was not the case for their control sample of inactive galaxies. They also showed that such young populations are not necessarily indicative of ongoing star formation. This is along the lines of the evolutionary sequence of Boisson et al. (2000, 2004). Based on the similar nuclear star formation histories of the Seyfert 2s and the star forming control sample galaxies, Burtscher et al. (2021) further speculate that the starbursts probably turned into Seyferts for some fraction of their time.
In the present paper, in the IR, we find that indeed star-forming galaxies may turn into Seyfert 1s, while star formation in Seyfert 2s, if any, appears outside the nuclear region. We note that NGC 2992, for example, is known to change from type 1 to type 2.